A sample of 0.53g of anhydrous CoCl2 is dissolved in 25.0 mL of water. Predict the absorbance at 510 nm for this solution, determined using the procedure described in this experiment.

Assume that ? = 3.15 for this problem.
Answer #1

no of moles of CoCl2 = 0.53/ 130 = 0.00408 moles

concentration = no of moles / volume = 0.00408/ 0.0250 = 0.1632 M

Absorbance =  ?l c = 3.15 x 1 x 0.1632 = 0.514